# Light-Weight Fault Tolerant Attention for Large Language Model Training

### v2

@@ Abstract.
Large Language Models (LLMs) have demonstrated remarkable performance in various natural language processing tasks. However, the training of these models is computationally intensive and susceptible to faults, particularly in the attention mechanism, which is a critical component of transformer-based LLMs. In this paper, we investigate the impact of faults on LLM training, focusing on INF, NaN, and near-INF values in the computation results with systematic fault injection experiments. We observe the propagation patterns of these errors, which can trigger non-trainable states in the model and disrupt training, forcing the procedure to load from checkpoints. To mitigate the impact of these faults, we propose ATTNChecker, the first Algorithm-Based Fault Tolerance (ABFT) technique tailored for the attention mechanism in LLMs. ATTNChecker is designed based on fault propagation patterns of LLM and incorporates performance optimization to adapt to both system reliability and model vulnerability while providing lightweight protection for fast LLM training. Evaluations on four LLMs show that ATTNChecker on average incurs on average 7% overhead on training while detecting and correcting all extreme errors. Compared with the state-of-the-art checkpoint/restore approach, ATTNChecker reduces recovery overhead by up to 49$\times$.
Fault toleration, ABFT, Attention mechanism, LLMS, Matrix multiplication
@@ 1. Introduction
Large Language Models (LLMs) commonly consume significant resources for training and fine-tuning. With the model consisting of more parameters and complexity, managing and updating this vast number of parameters during the training process demands substantial computational power. In addition, LLMs are trained on extensive collections of data, also consuming significant computing resources to finish the training steps. As a result, training an LLM is an extremely time and resource-intensive process, often taking weeks or even months to reach the desired accuracy. Most LLMs therefore are accelerated by high-performance systems, featuring a massive number of GPUs, dedicated interconnects, and so on. For example, training GPT-3 (Brown, 2020), reportedly took thousands of GPUs. The MegaScale project (Jiang et al., 2024) indicates the use of 10,000 GPUs to meet the demand of such large-scale training. Smaller models like BERT (Devlin et al., 2018), which has around 110 million parameters for its base version, might require fewer resources but are still GPU-intensive tasks.
@@
As systems scale up and GPUs are extensively utilized, it is risky to experience system-wise transient hardware faults, a phenomenon well-documented in the field report of large systems (Nie et al., 2016; Gonçalves de Oliveira et al., 2016; Martino et al., 2015; Dixit et al., 2022). Training LLMs with HPC resources needs to overcome the soft errors. The growing size and complexity of LLMs make it imperative to mitigate the impact of soft errors during training. Previous studies (Zhang et al., 2019) have explored the resilience of machine learning model training against minor data deviations caused by soft errors, concluding that such deviations often result in benign outcomes, with models typically converging successfully despite these errors. A more recent investigation (He et al., 2023) provides a comprehensive analysis through systematic fault injection experiments conducted during model training. This study is particularly interesting as it offers insights into the implications of specific faults, notably Infinity (INF) and Not-a-Number (NaN) errors, in the context of LLM training. Unlike normal error, LLM training typically cannot recover from INF or NaN error via additional training cycles.
@@
The best practice to handle the occurrence of INF and NaN during the training is for the model to roll back to a checkpoint where no INF and NaN are present and retrain the model from the stored parameters (Wang et al., 2023). While effective, retraining is considered a costly recovery method for LLM training in terms of time overhead, resource occupation and expense in dollars (Wu et al., 2023; Wan et al., 2024; Rajbhandari et al., 2020). Moreover, since INF and NaN can propagate during the computation, it might be necessary for the model to roll back to an earlier checkpoint that is steps away, which further exacerbates the repair cost.
@@
To overcome the outcome of INF and NAN on the model training and resolve the inefficiency introduced by the current solution, we introduce an innovative algorithm-based fault tolerance technique (a.k.a ABFT) optimized for LLM training. The proposed technique, denoted as ATTNChecker, is designed to efficiently and effectively identify and rectify abnormal values in real time during the training process. The primary objective of ATTNChecker is to develop a mechanism that is both lightweight and specifically tailored to the architectural needs of LLMs. ATTNChecker offers two advantages: First, the ability to detect anomalies where they initially occur prevents the further spread of errors through the system, thereby maintaining the integrity of the training process. Second, ABFT’s capacity for immediate correction enables the system to recalibrate any infinite (INF), undefined (NaN) or other types of errors that can eventually lead to INF or NAN (i.e. near-infinite (near-INF) values) through a fast correction process. This not only relaxes the need for extensive rollback procedures but also significantly reduces the dependency on checkpoint-based recovery methods, thereby streamlining the training process and enhancing the overall training efficiency.
@@
ATTNChecker aims to address the above challenges with the following assumptions. The attention mechanisms, particularly self-attention in Transformer-based LLMs, are among the most computationally demanding components in these models. This high demand stems from the quadratic increase in computational complexity with the length of input sequences, intensive matrix multiplication operations, and the large number of parameters within the attention layers. Furthermore, the attention mechanism’s frequent and intensive memory access can heavily occupy the system’s memory bandwidth. These factors combined make the attention mechanism a central focus not only in the purpose of optimizing the efficiency of LLMs, but also in improving the overall fault tolerance of LLM training.
@@
    We design the first ABFT that can effectively handle INF, NaN, and near-INF errors - Extreme Error Correcting ABFT (EEC-ABFT). EEC-ABFT is also the first ABFT that is highly optimized to handle errors in various circumstances including propagated errors, unpredictable patterns, and mixed error types.
@@
    Based on EEC-ABFT, we build the first comprehensive soft error protection approach for attention mechanism- ATTNChecker. ATTNChecker is specially tailored for protecting all major operations in attention and optimized for adaptive system reliability. As ATTNChecker is completely transparent to models, any existing LLMs can gain reliability improvement by integrating our ATTNChecker with minimum modification efforts, and in complementary with existing fault tolerance mechanisms.
@@
    We integrate ATTNChecker into the PyTorch framework. The evaluation shows that ATTNChecker on average incurs on average 7% overhead on training while reaching 100% detection and correction rate across all extreme errors. Our performance estimation shows that ATTNChecker reduces recovery overhead by up to 49×49$\times$49 × compared with the state-of-the-art recovery technique.
@@ 2.1. Attention in LLMs
The attention mechanism is the core building block in Large Language Models (LLMs). It maps a query and a set of key-value pairs to the result of a weighted sum, effectively providing a contextually relevant representation based on the input query and the key-value pairs (Vaswani et al., 2017). Query, Key and Value are matrices, each representing instance vectors. The Self-Attention could be described as:
@@
Based on Self-Attention the multi-head attention(Vaswani et al., 2017) is implemented by linearly projecting Q, K and V multiple times, then each head performs self-attention based on its own Q, K and V. The results of all heads will be combined. Multi-Head Attention is described as:
@@ 2.2. Soft Errors causing INF, NaN, or near-INF Values
According to IEEE754, the binary representation of floating point numbers includes 3 basic parts: sign, exponent and mantissa. Depending on where bit-flips occur they can affect the original value differently. Two specific types of floating-point errors can be caused by bit-flips: INF (i.e. infinity) and NaN (i.e. Not-a-Number). Although not classified as an exception by convention, extremely large values (near-INF) resulting from bit-flips in the exponent can also negatively impact computation. There are no specific rules for identifying near-INF, we generally consider values that are larger than a threshold as near-INF, which can vary for different computations. INF, NaN, and near-INF could cause several exceptions during the computation, as they can legally propagate through calculations e.g., INF times any number equals INF. In addition, one type of exception can transit to another type e.g., A near-INF number operating on another near-INF may lead to INF.
@@ V FI - - 1C-M 2D-M
Prior studies (He et al., 2023) have shown that a small perturbation affecting the LLM training data is likely to be mitigated over the model convergence process and the final model’s accuracy would be pertained under such errors. However, in terms of floating point values that belong to special categories like INF, NaN, and near-INF, prior studies have demonstrated that the LLM models would be seriously compromised under such errors, leading to unacceptable outcomes. Moreover, a near-INF number is also problematic as it can either accumulate and result in INF/NaN, or disrupt the gradient distribution to reduce the convergence probability.
To understand the different impacts of soft errors on LLM model training process, we conduct a fault injection and error propagation study on four representative LLMs: Bert, GPT-2, GPT-Neo, and Roberta. We specifically focus on studying the impact of faults that lead to INF, NaN, and near-INF, since those lead to the most drastic value change that may greatly impact model training. The goal of our study is to understand (1) how an error propagates across operations in attention modules; (2) How do such errors impact the training state? We hypothesize that the answers to these questions would guide the design of efficient and effective ABFT approaches for attention modules.
We select the compute-intensive GEMMs in the attention mechanism as the fault injection site since they consume most of the computation cycles for the entire training process. (Table 2 shows the workload ratios of GEMM in attention mechanism in four LLMs (Hoffmann et al., 2022).) In particular, we inject one fault (i.e., 0D pattern) into the result matrix of each GEMM to simulate a fault that occurred during the computation. Then, we carefully examine the errors in matrices involved in the following operations.
@@ 3.2. Insight for Designing Fault Tolerance
Segmented Protection: As error propagations could easily lead to a catastrophic result, operations in attention mechanism need to be segmented with fault detection at the operation boundary to confine the error. E.g., if a 2D pattern is beyond recoverable, then error detection/correction needs to be done for $AS$, $CL$, and $O$ to minimize the probability of the occurrence of a 2D pattern.
Handling (Non)deterministic Patterns: For operations with deterministic error patterns, efficient fault tolerance can be implemented as the deterministic pattern is bounded with no additional pattern detection required. For instance, if we use segmented fault tolerance, then $O$ will have only deterministic 1R patterns (0D pattern can be handled the same way as 1R). However, for operations with nondeterministic error patterns, fault tolerance needs to be adaptive to different potential patterns. For example, the error pattern of $AS$ can be either 1R or 1C depending on the origin of the error.
@@ 4. ABFT for Attention Mechanism
In this section, we propose to build the first ABFT for the attention mechanism, a key building block of many LLMs. This section begins with the introduction of our fault model, which defines the types of faults ATTNChecker is interested in. Then, we introduce ABFT with an extreme error correcting algorithm (EEC-ABFT) against INF, NaN, and near-INF errors and various types of propagated errors in attention. We then use EEC-ABFT to build a systematic protection for the attention mechanism denoted as ATTNChecker. In addition to providing comprehensive protection, ATTNChecker also makes ABFT detection adaptable to system reliability and model vulnerability. Finally, we build highly-optimized GPU kernels for achieving lightweight ABFT protection for the attention mechanism. Our goal is to maximize the coverage of attention mechanism against soft errors such that the reliability of LLM training can be greatly improved while the training performance is well-maintained.
@@ 4.1. Fault Model
The occurrence of INF, NaN, and near-INF can originate from a diverse array of sources: (1) Input-Related Faults: Certain inputs to the model can trigger floating-point exceptions during numerical operations, leading to INF/NaN values. (2) Hardware Faults: bit-flips induced by radiation or manufacturing defects can affect the correctness of computations, generating extreme values. Note that memory is usually ECC-protected in HPC systems hence a memory error is beyond the scope of our paper. (3) Inappropriate Hyperparameters: The use of unsuitable hyperparameters, like an excessively large learning rate, as indicated in prior research (Xu et al., 2021), can cause a rapid increase in the loss function (known as loss explosion), resulting in INF/NaN values
Input-related faults, likely caused by certain input data, cannot be effectively addressed using ABFT. This is because repeating the same computations during training would reproduce these exceptions. Conversely, issues related to loss explosion, often a consequence of inappropriate hyperparameters, can be dynamically mitigated through parameter adjustments. Therefore, the focus of the paper is to address the impact of transient hardware faults that lead to soft error caused INF, NaN, or near-INF in the attention mechanism. These hardware-related errors are particularly problematic as they not only cause severe disruptions in the training process but also entail substantial overhead for correction in existing training frameworks. As attention mechanism is one the most important building blocks of many LLMs, this work targets soft errors that occur during the execution of attention mechanism.
@@ 4.2. Extreme Errors Correcting ABFT
Case 1: $\delta_{1}<INF$ indicates all error(s) are less than INF. Then we count the number of elements that are greater than a threshold $T_{near-INF}$ for identifying near-INF. If such a number $K$ is one, then it indicates that the error is not propagated. Next, we precisely locate the error. Normally, one can calculate the error index using $\delta_{2}/\delta_{1}$, but we must be careful since the weighted checksum of a vector containing near-INF elements can cause overflow in $\delta_{2}$. So, EEC-ABFT checks the value of $\delta_{2}$ before locating the error. If $\delta_{2}<INF$, we proceed with the regular locating logic. Otherwise, we search for the largest value in $v$ to find the error index. Finally, EEC-ABFT corrects the error given the location. Special attention is worth paying to error correction for near-INF cases, since a round-off error may cause corrupted value to be wrongly corrected. In this case, $\delta_{1}$ will absorb smaller $v_{j}$ in the correction process leading to incorrect recovery. To avoid such precision-related error, we define a threshold $T_{correct}$ to determine which error correcting strategy to use. Empirically, we use $T_{near-INF}=10^{10}$ and $T_{correct}=10^{5}$.
@@
Case 4: For all previous three cases, we check if there is an error propagation in $v$ by counting the number of errors. If it is greater than one, a 1D error propagation is identified, then EEC-ABFT will abort the current correction process as we need to initiate special procedure for error propagation. (will be discussed in Section 4.3).

### v3

# ATTNChecker: Highly-Optimized Fault Tolerant Attention for Large Language Model Training

@@ Abstract.
Large Language Models (LLMs) have demonstrated remarkable performance in various natural language processing tasks. However, the training of these models is computationally intensive and susceptible to faults, particularly in the attention mechanism, which is a critical component of transformer-based LLMs. In this paper, we investigate the impact of faults on LLM training, focusing on INF, NaN, and near-INF values in the computation results with systematic fault injection experiments. We observe the propagation patterns of these errors, which can trigger non-trainable states in the model and disrupt training, forcing the procedure to load from checkpoints. To mitigate the impact of these faults, we propose ATTNChecker, the first Algorithm-Based Fault Tolerance (ABFT) technique tailored for the attention mechanism in LLMs. ATTNChecker is designed based on fault propagation patterns of LLM and incorporates performance optimization to adapt to both system reliability and model vulnerability while providing lightweight protection for fast LLM training. Evaluations on four LLMs show that ATTNChecker incurs on average 7% overhead on training while detecting and correcting all extreme errors. Compared with the state-of-the-art checkpoint/restore approach, ATTNChecker reduces recovery overhead by up to 49$\times$.
Algorithm-Based Fault Tolerance, Attention Mechanism, Large Language Models, Matrix Multiplication
@@
journalyear: 2025copyright: ccconference: The 30th ACM SIGPLAN Annual Symposium on Principles and Practice of Parallel Programming; March 1–5, 2025; Las Vegas, NV, USAbooktitle: The 30th ACM SIGPLAN Annual Symposium on Principles and Practice of Parallel Programming (PPoPP ’25), March 1–5, 2025, Las Vegas, NV, USAdoi: 10.1145/3710848.3710870isbn: 979-8-4007-1443-6/25/03ccs: Computing methodologies Machine learningccs: Computer systems organization Parallel architecturesccs: Computer systems organization Dependable and fault-tolerant systems and networks
@@ 1. Introduction
Large Language Models (LLMs) commonly consume significant resources for training and fine-tuning. With the model consisting of more parameters and complexity, managing and updating this vast number of parameters during the training process demands substantial computational power. In addition, LLMs are trained on extensive collections of data, also consuming significant computing resources to finish the training steps. As a result, training an LLM is an extremely time and resource-intensive process, often taking weeks or even months to reach the desired accuracy. Most LLMs, therefore, are accelerated by high-performance systems featuring a massive number of GPUs, dedicated interconnects, and so on. For example, training GPT-3 (Brown, 2020), reportedly took thousands of GPUs. The MegaScale project (Jiang et al., 2024) indicates using 10,000 GPUs to meet the demand of such large-scale training. Smaller models like BERT (Devlin et al., 2018), which has around 110 million parameters for its base version, might require fewer resources but are still GPU-intensive tasks.
@@
As systems scale up and GPUs are extensively utilized, it is risky to experience system-wise transient hardware faults, a phenomenon well-documented in the field report of large systems (Nie et al., 2016; Gonçalves de Oliveira et al., 2016; Martino et al., 2015; Dixit et al., 2022). Training LLMs with HPC resources needs to overcome the soft errors. The growing size and complexity of LLMs make it imperative to mitigate the impact of soft errors during training. Previous studies (Zhang et al., 2019) have explored the resilience of machine learning model training against minor data deviations caused by soft errors, concluding that such deviations often result in benign outcomes, with models typically converging successfully despite these errors. A more recent investigation (He et al., 2023) provides a comprehensive analysis through systematic fault injection experiments conducted during model training. This study is particularly interesting as it offers insights into the implications of specific faults, notably Infinity (INF) and Not-a-Number (NaN) errors, in the context of LLM training. Unlike normal errors, LLM training typically cannot recover from INF or NaN errors via additional training cycles.
@@
The best practice to handle the occurrence of INF and NaN during the training is for the model to roll back to a checkpoint where no INF and NaN are present and retrain the model from the stored parameters (Wang et al., 2023). While effective, retraining is considered a costly recovery method for LLM training in terms of time overhead, resource occupation, and expense in dollars (Wu et al., 2023; Wan et al., 2024; Rajbhandari et al., 2020). Moreover, since INF and NaN can propagate during the computation, it might be necessary for the model to roll back to an earlier checkpoint that is steps away, which further exacerbates the repair cost.
@@
To overcome the outcome of INF and NAN on the model training and resolve the inefficiency introduced by the current solution, we introduce an innovative algorithm-based fault tolerance technique (a.k.a ABFT) optimized for LLM training. The proposed technique, denoted as ATTNChecker, is designed to efficiently and effectively identify and rectify abnormal values in real-time during training. The primary objective of ATTNChecker is to develop a mechanism that is both lightweight and specifically tailored to the architectural needs of LLMs. ATTNChecker offers two advantages: First, the ability to detect anomalies where they initially occur prevents the further spread of errors through the system, thereby maintaining the integrity of the training process. Second, ABFT’s capacity for immediate correction enables the system to recalibrate any infinite (INF), undefined (NaN), or other types of errors that can eventually lead to INF or NAN (i.e., near-infinite (near-INF) values) through the fast correction process. This not only relaxes the need for extensive rollback procedures but also significantly reduces the dependency on checkpoint-based recovery methods, thereby streamlining the training process and enhancing the overall training efficiency.
@@
ATTNChecker aims to address the above challenges with the following assumptions. The attention mechanisms, particularly self-attention in Transformer-based LLMs, are among the most computationally demanding components in these models. This high demand stems from the quadratic increase in computational complexity with the length of input sequences, intensive matrix multiplication operations, and the large number of parameters within the attention layers. Furthermore, the attention mechanism’s frequent and intensive memory access can heavily occupy the system’s memory bandwidth. These factors combined make the attention mechanism a central focus not only for the purpose of optimizing the efficiency of LLMs but also for improving the overall fault tolerance of LLM training.
@@
    We design the first ABFT that can effectively handle INF, NaN, and near-INF errors - Extreme Error Correcting ABFT (EEC-ABFT). EEC-ABFT is also the first ABFT that is highly optimized to handle errors in various circumstances, including propagated errors, unpredictable patterns, and mixed error types.
@@
    Based on EEC-ABFT, we build the first comprehensive soft error protection approach for attention mechanism- ATTNChecker. ATTNChecker is specially tailored for protecting all major operations in attention and optimized for adaptive system reliability. As ATTNChecker is completely transparent to models, any existing LLMs can improve reliability by integrating our ATTNChecker with minimum modification efforts and complementary with existing fault tolerance mechanisms.
@@
    We integrate ATTNChecker into the PyTorch framework. The evaluation shows that ATTNChecker incurs on average 7% overhead on training while reaching 100% detection and correction rate across all extreme errors. Our performance estimation shows that ATTNChecker reduces recovery overhead by up to 49×49$\times$49 × compared with the state-of-the-art recovery technique.
@@ 2.1. Attention in LLMs
The attention mechanism is the core building block in Large Language Models (LLMs). It maps a query and a set of key-value pairs to the result of a weighted sum, effectively providing a contextually relevant representation based on the input query and the key-value pairs (Vaswani et al., 2017). Query, Key, and Value are matrices, each representing instance vectors. The Self-attention could be described as:
@@
Based on self-attention, the multi-head attention(Vaswani et al., 2017) is implemented by linearly projecting Q, K, and V multiple times, then each head performs self-attention based on its own Q, K, and V. The results of all heads will be combined. Multi-Head Attention is described as:
@@ 2.2. Soft Errors causing INF, NaN, or near-INF Values
According to IEEE754, the binary representation of floating point numbers includes 3 basic parts: sign, exponent, and mantissa. Depending on where bit-flips occur they can affect the original value differently. Two specific types of floating-point errors can be caused by bit-flips: INF (i.e., infinity) and NaN (i.e., Not-a-Number). Although not classified as an exception by convention, extremely large values (near-INF) resulting from bit-flips in the exponent can also negatively impact computation. There are no specific rules for identifying near-INF, we generally consider values that are larger than a threshold as near-INF, which can vary for different computations. INF, NaN, and near-INF could cause several exceptions during the computation, as they can legally propagate through calculations e.g., INF times any number equals INF. In addition, one type of exception can transit to another type e.g., A near-INF number operating on another near-INF may lead to INF.
@@ 2.3. ABFT for Matrix Multiplication
Table 1. Table of Notation
Notation Description
$Q$ Query
$K$ Key
$V$ Value
$AS$ Attention Score
$AP$ Attention Probability
$CL$ Context Layer
$O$ Output
@@ V FI - - 1C-M 2D-M
Prior studies (He et al., 2023) have shown that a small perturbation affecting the LLM training data is likely to be mitigated over the model convergence process, and the final model’s accuracy would be pertained under such errors. However, in terms of floating point values that belong to special categories like INF, NaN, and near-INF, prior studies have demonstrated that the LLM models would be seriously compromised under such errors, leading to unacceptable outcomes. Moreover, a near-INF number is also problematic as it can either accumulate and result in INF/NaN or disrupt the gradient distribution to reduce the convergence probability.
To understand the different impacts of soft errors on the LLM model training process, we conduct a fault injection and error propagation study on four representative LLMs: Bert, GPT-2, GPT-Neo, and Roberta. We specifically focus on studying the impact of faults that lead to INF, NaN, and near-INF since those lead to the most drastic value change that may greatly impact model training. The goal of our study is to understand (1) how an error propagates across operations in attention modules; (2) How do such errors impact the training state? We hypothesize that the answers to these questions would guide the design of efficient and effective ABFT approaches for attention modules.
We select the compute-intensive GEMMs in the attention mechanism as the fault injection site since they consume most of the computation cycles for the entire training process. (Table 3 shows the workload ratios of GEMM in attention mechanism in four LLMs (Hoffmann et al., 2022).) In particular, we inject one fault (i.e., 0D pattern) into the result matrix of each GEMM to simulate a fault that occurred during the computation. Then, we carefully examine the errors in matrices involved in the following operations. The notations are presented in Table 1
@@ 3.2. Insight for Designing Fault Tolerance
Segmented Protection: As error propagations could easily lead to a catastrophic result, operations in attention mechanism need to be segmented with fault detection at the operation boundary to confine the error. For example, if a 2D pattern is beyond recoverable, then error detection/correction needs to be done for $AS$, $CL$, and $O$ to minimize the probability of the occurrence of a 2D pattern.
Handling (Non) deterministic Patterns: For operations with deterministic error patterns, efficient fault tolerance can be implemented as the deterministic pattern is bounded with no additional pattern detection required. For instance, if we use segmented fault tolerance, then $O$ will have only deterministic 1R patterns (0D pattern can be handled the same way as 1R). However, for operations with nondeterministic error patterns, fault tolerance needs to be adaptive to different potential patterns. For example, the error pattern of $AS$ can be either 1R or 1C, depending on the origin of the error.
@@ 4. ABFT for Attention Mechanism
In this section, we propose to build the first ABFT for the attention mechanism, a key building block of many LLMs. This section introduces our fault model, which defines the types of faults ATTNChecker is interested in. Then, we introduce ABFT with an extreme error correcting algorithm (EEC-ABFT) against INF, NaN, and near-INF errors and various types of propagated errors in attention. We then use EEC-ABFT to build a systematic protection for the attention mechanism denoted as ATTNChecker. In addition to providing comprehensive protection, ATTNChecker also makes ABFT detection adaptable to system reliability and model vulnerability. Finally, we build highly-optimized GPU kernels for achieving lightweight ABFT protection for the attention mechanism. Our goal is to maximize the coverage of attention mechanism against soft errors such that the reliability of LLM training can be greatly improved while the training performance is well-maintained.
@@ 4.1. Fault Model
The occurrence of INF, NaN, and near-INF can originate from diverse sources: (1) Input-Related Faults: Certain inputs to the model can trigger floating-point exceptions during numerical operations, leading to INF/NaN values. (2) Hardware Faults: Bit-flips induced by radiation or manufacturing defects can affect the correctness of computations, generating extreme values. Note that memory is usually ECC-protected in HPC systems; hence, a memory error is beyond the scope of our paper. (3) Inappropriate Hyperparameters: The use of unsuitable hyperparameters, like an excessively large learning rate, as indicated in prior research (Xu et al., 2021), can cause a rapid increase in the loss function (known as loss explosion), resulting in INF/NaN values.
Input-related faults, likely caused by certain input data, cannot be effectively addressed using ABFT. This is because repeating the same computations during training would reproduce these exceptions. Conversely, issues related to loss explosion, often a consequence of inappropriate hyperparameters, can be dynamically mitigated through parameter adjustments. Therefore, the focus of the paper is to address the impact of transient hardware faults that lead to soft errors caused INF, NaN, or near-INF in the attention mechanism. These hardware-related errors are particularly problematic as they not only cause severe disruptions in the training process but also entail substantial overhead for correction in existing training frameworks. As attention mechanism is one the most important building blocks of many LLMs, this work targets soft errors that occur during the execution of attention mechanism.
@@ 4.2. Extreme Errors Correcting ABFT
Case 1: $\delta_{1}<INF$ indicates all error(s) are less than INF. Then we count the number of elements that are greater than a threshold $T_{near-INF}$ for identifying near-INF. If such a number $K$ is one, then it indicates that the error is not propagated. Next, we precisely locate the error. Normally, one can calculate the error index using $\delta_{2}/\delta_{1}$, but we must be careful since the weighted checksum of a vector containing near-INF elements can cause overflow in $\delta_{2}$. So, EEC-ABFT checks the value of $\delta_{2}$ before locating the error. If $\delta_{2}<INF$, we proceed with the regular locating logic. Otherwise, we search for the largest value in $v$ to find the error index. Finally, EEC-ABFT corrects the error given the location. Special attention is worth paying to error correction for near-INF cases since a round-off error may cause a corrupted value to be wrongly corrected. In this case, $\delta_{1}$ will absorb smaller $v_{j}$ in the correction process leading to incorrect recovery. To avoid such precision-related errors, we define a threshold $T_{correct}$ to determine which error-correcting strategy to use. Empirically, we use $T_{near-INF}=10^{10}$ and $T_{correct}=10^{5}$.
@@
Case 4: For all the previous three cases, we check if there is an error propagation in $v$ by counting the number of errors. If it is greater than one, a 1D error propagation is identified, then EEC-ABFT will abort the current correction process as we need to initiate a special procedure for error propagation. (will be discussed in Section 4.3).